OpenStudy (anonymous):

Use quadratic formula to solve? 2x^2+5x-6=0

OpenStudy (accessdenied):

\[ x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \] a = 2, b = 5, c = -6

OpenStudy (anonymous):

\[So...-5\pm \sqrt{(5)^{2}}-4(2)(-6)\]

OpenStudy (anonymous):

/2(2)

OpenStudy (accessdenied):

yes, although the square root should also include the - 4(2)(-6).

OpenStudy (anonymous):

So, \[-5\sqrt{25-48}/4\]

OpenStudy (accessdenied):

-5 +- sqrt(25 - 4(2)(-6) ) ---------------------- 4 try rechecking that - 4(2)(-6).. There are two negatives being multiplied together (-4, -6) so you should be getting a positive.

OpenStudy (anonymous):

Dang it...yeah, so -5*-73/4

OpenStudy (anonymous):

Is it 25 + 48?

OpenStudy (accessdenied):

yes im not sure why you are multiplying the -5 to the root, it should be a plus-minus sign

OpenStudy (anonymous):

That was a typo, I meant to put +-

OpenStudy (accessdenied):

ah, okay. but yes, you would be correct with 25 + 48.

OpenStudy (anonymous):

So do you get x = -3.2679?

OpenStudy (anonymous):

Wait...just a sec

OpenStudy (anonymous):

-2.86399

OpenStudy (accessdenied):

i got... -5 + sqrt(73) -5 - sqrt(73) -------------, ------------- 4 4

OpenStudy (anonymous):

Oh, just leave it like that?

OpenStudy (accessdenied):

yeah, it doesn't ask for approximations, so it would be sufficient to keep it in that form

OpenStudy (anonymous):

OMG...thank you. There's so many steps to mess up on. Thanks for your time

OpenStudy (accessdenied):

Yeah, the tricky part (aside from memorizing the formula itself) is usually just getting all the arithmetic correct. :P